# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 2. Preliminary results
Let $0<b\leq\frac{n}{2}$ and $u\in E\setminus\{0\}$ be such that $\|\Delta u\|^{a}_{\frac{n}{2}}+\|\Delta u\|^{b}_{p}\leq 1$. Define
@@
If $\frac{1}{2^{\frac{n-2p}{p}}}<\theta<1$, then we define
@@
$\leq\lambda^{n-\gamma}\text{ATSC}(\theta^{\frac{n}{n-2}}\beta(n,2),\gamma)\leq\bigg{(}\frac{(1-\theta^{a})^{\frac{1}{b}}}{\theta}\bigg{)}^{p^{\ast}(1-\frac{\gamma}{n})}\frac{C(n,\gamma,p)}{(1-\theta^{\frac{n}{2}})^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}$
$\leq\frac{(1-\theta^{a})^{\frac{p^{\ast}}{b}(1-\frac{\gamma}{n})}}{(1-\theta^{\frac{n}{2}})^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}C(n,\gamma,p)\leq C(n,\gamma,p,a,b),$
thanks to the fact that $b\leq\frac{n}{2}$, where $C(n,\gamma,p,a,b)>0$ is a constant depending only on $n,\gamma,p,a$ and $b$. If $0<\theta\leq\frac{1}{2^{\frac{n-2p}{p}}}$, then we consider
@@
$\leq 2^{n-\gamma}\frac{C(n,\gamma,p)}{\Big{(}1-2^{-\frac{n^{2}}{2p^{\ast}}}\Big{)}^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}\leq C(n,\gamma,p),$
where $C(n,\gamma,p)>0$ is a constant depending only on $n,~{}\gamma$ and $p$. In conclusion, we deduce that $\text{ATC}_{a,b}(\gamma)<+\infty$ for $b\leq\frac{n}{2}$. Next, to prove the sharpness of $\beta(n,2)$, inspired by Theorem 1.5, we assume that
@@
where $v$ is defined as in (4.3). Let us choose
@@
Notice that $\|\Delta\bar{w}\|^{a}_{\frac{n}{2}}+\|\Delta\bar{w}\|^{b}_{p}=1$. Hence, by choosing $\ell>\beta(n,2)$, we obtain
@@
$=\frac{\omega_{n-1}}{n-\gamma}~{}\operatorname{\exp}\Bigg{[}n(1-\frac{\gamma}{n}\Big{)}\log\frac{1}{r}\Bigg{(}\frac{\ell}{\beta(n,2)}\frac{\delta_{r}^{\frac{n}{n-2}}}{\Big{(}1+2\varepsilon\Big{(}\|f^{\prime}\|_{\infty}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Big{)}^{\frac{n}{2}}\Big{)}^{\frac{2}{n-2}}}-1\Bigg{)}\Bigg{]}$
@@
$\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma)\leq\text{ATC}_{a,b}(\gamma).$ (4.4)
@@
Define
@@
Consequently, we obtain from the definition of $\text{ATSC}(\ell,\gamma)$ that
@@
$\leq\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma).$
@@
$\text{ATC}_{a,b}(\gamma)\leq\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma).$ (4.5)
In view of (4.4) and (4.5), we deduce that (1.15) holds.
To complete the proof, we claim that $\text{ATC}_{a,b}(\gamma)<+\infty$ implies $b\leq\frac{n}{2}$. Indeed, if not, let there exists some $b>\frac{n}{2}$ such that $\text{ATC}_{a,b}(\gamma)<+\infty$. Thus, one can obtain from (1.15) that
$\limsup_{\ell\to\beta(n,2)^{-}}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma)<+\infty.$ (4.6)
@@
$\limsup_{\ell\to\beta(n,2)^{-}}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma)>0.$ (4.7)
@@
[…]
which is a contradiction, thanks to $b>\frac{n}{2}$ and hence the claim. This finishes the proof. ∎
5. Proof of Theorem 1.7
This section concerns about the attainability of the best constant in the embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$. For this purpose, we first need to prove Theorem 1.7.
@@
First, we claim that the continuous embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ holds for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$. Indeed, let $v\in E$ and $\eta,~{}\eta^{\prime}>1$ be such that $\frac{1}{\eta}+\frac{1}{\eta^{\prime}}=1$ satisfying $\gamma\eta^{\prime}<n$. By using Hölder’s inequality and Lemma 2.5, one has
$\int_{\mathbb{R}^{n}}\frac{|v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{B_{1}}\frac{|v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x+\int_{B^{c}_{1}}\frac{|v|^{\varrho}}{|x|^{\gamma}}~{}\mathrm{d}x$
$\leq\bigg{(}\int_{B_{1}}|v|^{\varrho\eta}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta}}\bigg{(}\int_{B_{1}}\frac{1}{|x|^{\gamma\eta^{\prime}}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta^{\prime}}}+\int_{\mathbb{R}^{n}}|v|^{\varrho}\,\mathrm{d}x\lesssim\|v\|^{\varrho},$
which conclude the claim. It remains to check the compactness of the above embedding. To prove this, take a bounded sequence $\{v_{k}\}_{k}\subset E$, then one can find a subsequence $\{v_{k_{j}}\}_{j}$ and $v$ in $E$ satisfying
$v_{k_{j}}\rightharpoonup v~{}\text{in}~{}E,~{}v_{k_{j}}\to v~{}\text{in}~{}L^{\theta}(B_{R})\quad\text{for any}~{}R>0~{}\text{and}~{}\theta\geq 1,~{}v_{k_{j}}\to v~{}\text{a.e. in}~{}\mathbb{R}^{n}~{}\text{as}~{}j\to\infty,$
we thank to Lemma 2.5. Due to the Egoroff’s theorem, we can notice that for any open ball $B_{R}$ and $\delta>0$, there exists $A_{\delta}\subset B_{R}$ with $|A_{\delta}|<\delta$ satisfying $v_{k_{j}}\to v$ uniformly in $B_{R}\setminus A_{\delta}$ as $j\to\infty$. Further, let us write
$\lim_{j\to\infty}\int_{\mathbb{R}^{n}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$ $=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$
@@
$\quad+\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}\setminus B_{R}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x:=J_{1}+J_{2}+J_{3}.$ (5.1)
Suppose that $1<t<\frac{n}{\gamma}$ and $1<r<\frac{n}{t\gamma}$ such that $\frac{1}{\tau}+\frac{1}{\tau^{\prime}}=1$ for any $\tau\in\{t,r\}$. Now by using Hölder’s inequality and Lemma 2.5, we have
@@
$J_{1}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (5.2)
Further, we also have
@@
$J_{2}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{B_{R}\setminus A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (5.3)
@@
$J_{3}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}\setminus B_{R}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (5.4)
In virtue of (5.2), (5.3) and (5.4), we obtain from (5) that $v_{k_{j}}\to v$ in $L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ as $j\to\infty$. This completes the proof. ∎
@@
$v_{k}\rightharpoonup v~{}\text{in}~{}E,~{}v_{k}\to v~{}\text{in}~{}L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x),~{}\forall~{}\varrho\geq p^{\ast}~{}\text{and}~{}v_{k}\to v~{}\text{a.e. in}~{}\mathbb{R}^{n}~{}\text{as}~{}k\to\infty,$
thanks to Theorem 1.7. This infers that $\|v\|_{\varrho,\gamma}=\lim_{k\to\infty}\|v_{k}\|_{\varrho,\gamma}=1$. Moreover, due to the weakly lower semicontinuity of the norm, we have $\|v\|\leq\liminf_{k\to\infty}\|v_{k}\|=S_{\varrho}\leq\|v\|$. It follows that the best constant $S_{\varrho}$ is achieved. This finishes the proof. ∎
6. Proof of Theorem 1.8
In this section, we shall use the mountain pass theorem, standard topological tools, singular Adams’ inequality, a new concentration-compactness principle due to Lions (see, for example, Theorem 1.3 and Theorem 1.4 ) and Theorem 1.7, respectively, to prove Theorem 1.8. Evidently, the problem ($\boldsymbol{S}$) has a variational structure. Indeed, since we are interested in studying ($\boldsymbol{S}$), we define the Euler–Lagrange variational energy functional $J\colon E\to\mathbb{R}$ by
$J(u):=\frac{1}{p}\|\Delta u\|^{p}_{p}+\frac{2}{n}\|\Delta u\|^{\frac{n}{2}}_{\frac{n}{2}}-\int_{\mathbb{R}^{n}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x,~{}\forall~{}u\in E.$ (6.1)
@@
$\langle{J^{\prime}(u),v}\rangle:=\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u\Delta v\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u\Delta v\,\mathrm{d}x-\int_{\mathbb{R}^{n}}\frac{g(x,u)}{|x|^{\gamma}}v\,\mathrm{d}x,~{}\forall~{}u,v\in E.$
Moreover, the critical points of $J$ are exactly weak solutions of ($\boldsymbol{S}$). To prove our existence results, we need a version of the mountain pass theorem of Rabinowitz [42], as stated below.
@@
$\mathcal{I}(u_{k})\to c\quad\text{and}\quad\sup_{\|\varphi\|_{X}=1}|\langle{\mathcal{I}^{\prime}(u_{k}),\varphi}\rangle|\to 0\quad\text{as}\quad k\to\infty$ (6.2)
@@
(The mountain pass theorem ) Let $X$ be a real Banach space and $I\in C^{1}(X,\mathbb{R})$. Suppose $I(0)=0$ and there hold
@@
    there exist two constants $\alpha,\rho>0$ such that $I(u)\geq\alpha$ for all $u\in X$ with $\|u\|=\rho$;
@@
    there exists $e\in X$ satisfying $\|e\|>\rho$ such that $I(e)<0$.
@@
$\Gamma:=\{\gamma\in C^{1}([0,1],X)\colon\gamma(0)=0,~{}\gamma(1)=e\}.$
@@
$c:=\inf_{\gamma\in\Gamma}\max_{t\in[0,1]}I(\gamma(t))\geq\alpha$
and there exists a $(PS)_{c}$ sequence $\{u_{k}\}_{k}$ for $I$ in $X$. Consequently, if $I$ satisfies the Palais-Smale condition, then $I$ possesses a critical value $c\geq\alpha$.
@@
Let the hypotheses $(g_{1})-(g_{4})$ be satisfied. Then the following results are true:
@@
Let $t,t^{\prime}>1$ be such that $\frac{1}{t}+\frac{1}{t^{\prime}}=1$. Further, we assume $\kappa>0$ such that $\alpha t^{\prime}\kappa^{\frac{n}{n-2}}\leq\beta_{\gamma,n}$ whenever $\|u\|\leq\kappa$. From (1.2), we obtain by using Hölder’s inequality, Theorem 1.7 and Corollary 2.6 that
@@
$\leq C(\zeta\|u\|^{\tau}+\|u\|^{q}),\quad\forall\quad\zeta>0,$
where we used Theorem 1.3 to obtain the constant $C>0$, which is given by

### v3

@@ 2. Preliminary results
6. Equivalence of critical and subcritical sharp singular Adams’ type inequalities: Proof of Theorem 1.4
The purpose of this section is to establish the relationship between the suprema for the critical and subcritical sharp singular Adams’ type inequalities. Initially, we prove the important lemma shown below.
Lemma 6.1.
Proof.
Let $u\in E$ be such that $\|\Delta u\|_{\frac{n}{2}}\leq 1$ and $\|\Delta u\|_{p}=1$. Define a new function $w$ as follows
$w(x)=\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{n}}u(\lambda x)\quad\text{with}\quad\lambda=\Bigg(\cfrac{(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}\Bigg)^{\frac{p}{b(n-2p)}}.$
By direct calculations, we have
$\|\Delta w\|^{a}_{\frac{n}{2}}=\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\big(\frac{n-2}{n}\big)a}\|\Delta u\|^{a}_{\frac{n}{2}}\leq\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\big(\frac{n-2}{n}\big)a}$
and
$\|\Delta w\|^{b}_{p}$ $=\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{(\frac{n-2}{n})b}\lambda^{-b\big(\frac{n-2p}{p}\big)}\|\Delta u\|^{b}_{p}=1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{(\frac{n-2}{n})a}.$
It follows that $\|\Delta w\|^{a}_{\frac{n}{2}}+\|\Delta w\|^{b}_{p}\leq 1$. Now, we deduce from the definition of $\text{ATC}_{a,b}(\gamma)$ that
$\leq\Bigg(\cfrac{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}\Bigg)^{\frac{p^{\ast}}{b}\big(1-\frac{\gamma}{n}\big)}\text{ATC}_{a,b}(\gamma).$
The above inequality ensures that the proof of Lemma 6.1 is finished at this point. ∎
Now, we are in a suitable position to prove Theorem 1.4 as follows.
@@
Let $0<b\leq\frac{n}{2}$ and $u\in E\setminus\{0\}$ be such that $\|\Delta u\|^{a}_{\frac{n}{2}}+\|\Delta u\|^{b}_{p}\leq 1$. Now, we define
@@
If $\cfrac{1}{2^{\frac{n-2p}{p}}}<\theta<1$, then we again define
@@
$\leq\lambda^{n-\gamma}\text{ATSC}(\theta^{\frac{n}{n-2}}\beta(n,2),\gamma)\leq\bigg(\frac{(1-\theta^{a})^{\frac{1}{b}}}{\theta}\bigg)^{p^{\ast}(1-\frac{\gamma}{n})}\frac{C(\gamma,p,n)}{(1-\theta^{\frac{n}{2}})^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}$
$\leq\frac{(1-\theta^{a})^{\frac{p^{\ast}}{b}(1-\frac{\gamma}{n})}}{(1-\theta^{\frac{n}{2}})^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}C(\gamma,p,n)\leq C(\gamma,p,n,a,b),$
thanks to the fact that $b\leq\frac{n}{2}$, where $C(\gamma,p,n,a,b)>0$ is a constant depending only on $\gamma,p,n,a$ and $b$. If $0<\theta\leq\cfrac{1}{2^{\frac{n-2p}{p}}}$, then we consider
@@
$\leq 2^{n-\gamma}\frac{C(\gamma,p,n)}{\Big(1-2^{-\frac{n^{2}}{2p^{\ast}}}\Big)^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}\leq C(\gamma,p,n),$
where $C(\gamma,p,n)>0$ is a constant depending only on $\gamma,n$ and $p$. Hence, we deduce that $\text{ATC}_{a,b}(\gamma)<+\infty$ for $b\leq\frac{n}{2}$. Next, to prove the sharpness of $\beta(n,2)$, inspired by Theorem 1.3, we assume that
@@
where $v$ is defined as in (5.3). Further, we define another function $\bar{w}$ as follows
@@
Notice that $\|\Delta\bar{w}\|^{a}_{\frac{n}{2}}+\|\Delta\bar{w}\|^{b}_{p}=1$. Moreover, by choosing $\ell>\beta(n,2)$, we obtain
@@
$=\frac{\omega_{n-1}}{n-\gamma}~\operatorname{\exp}\Bigg[n\Big(1-\frac{\gamma}{n}\Big)\log\frac{1}{r}\Bigg(\frac{\ell}{\beta(n,2)}\frac{\delta_{r}^{\frac{n}{n-2}}}{\Big(1+2\varepsilon\Big(\|f^{\prime}\|_{\infty}+O\Big(\Big(\log\frac{1}{r}\Big)^{-1}\Big)\Big)^{\frac{n}{2}}\Big)^{\frac{2}{n-2}}}-1\Bigg)\Bigg]$
@@
(6.1) $\sup_{\ell\in(0,\beta(n,2))}\Bigg(\cfrac{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}\Bigg)^{\frac{p^{\ast}}{b}\big(1-\frac{\gamma}{n}\big)}\text{ATSC}(\ell,\gamma)\leq\text{ATC}_{a,b}(\gamma).$
@@
For each $k\in\mathbb{N}$, we define the sequence $\{w_{k}\}_{k}$ as follows
@@
Hence, we obtain from the definition of $\text{ATSC}(\ell,\gamma)$ that
@@
$\leq\sup_{\ell\in(0,\beta(n,2))}\Bigg(\cfrac{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}\Bigg)^{\frac{p^{\ast}}{b}\big(1-\frac{\gamma}{n}\big)}\text{ATSC}(\ell,\gamma).$
@@
(6.2) $\text{ATC}_{a,b}(\gamma)\leq\sup_{\ell\in(0,\beta(n,2))}\Bigg(\cfrac{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}\Bigg)^{\frac{p^{\ast}}{b}\big(1-\frac{\gamma}{n}\big)}\text{ATSC}(\ell,\gamma).$
In view of (6.1) and (6.2), we deduce that (1.13) holds. To complete the proof, we claim that $\text{ATC}_{a,b}(\gamma)<+\infty$ implies $b\leq\frac{n}{2}$. Indeed, if not, let there exists some $b>\frac{n}{2}$ such that $\text{ATC}_{a,b}(\gamma)<+\infty$. Thus, one can obtain from (1.13) that
(6.3) $\limsup_{\ell\to\beta(n,2)^{-}}\Bigg(\cfrac{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}\Bigg)^{\frac{p^{\ast}}{b}\big(1-\frac{\gamma}{n}\big)}\text{ATSC}(\ell,\gamma)<+\infty.$
@@
(6.4) $\limsup_{\ell\to\beta(n,2)^{-}}\Bigg(\cfrac{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}\Bigg)^{\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}\text{ATSC}(\ell,\gamma)>0.$
@@
$\liminf_{\ell\to\beta(n,2)^{-}}\Bigg(\cfrac{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}\Bigg)^{\frac{p^{\ast}}{b}\big(1-\frac{\gamma}{n}\big)}\Bigg(\cfrac{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}\Bigg)^{\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}<+\infty,$
which is a contradiction, thanks to $b>\frac{n}{2}$ and thus, the claim is verified. This finishes the proof. ∎
7. The compact embedding result: Proof of Theorem 1.5
This section is concerned with the attainability of the best constant in the embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$. For this purpose, we first need to prove Theorem 1.5.
@@
First, we claim that the continuous embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ holds for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$. Indeed, let $v\in E$ and $\eta,~\eta^{\prime}>1$ be such that $\frac{1}{\eta}+\frac{1}{\eta^{\prime}}=1$ satisfying $\gamma\eta^{\prime}<n$. By using the Hölder’s inequality and Lemma 2.5, one has
$\int_{\mathbb{R}^{n}}\frac{|v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$ $=\bigg(\int_{B_{1}}+\int_{B^{c}_{1}}\bigg)\frac{|v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x\leq\bigg(\int_{B_{1}}|v|^{\varrho\eta}\,\mathrm{d}x\bigg)^{\frac{1}{\eta}}\bigg(\int_{B_{1}}\frac{1}{|x|^{\gamma\eta^{\prime}}}\,\mathrm{d}x\bigg)^{\frac{1}{\eta^{\prime}}}+\int_{\mathbb{R}^{n}}|v|^{\varrho}\,\mathrm{d}x\lesssim\|v\|^{\varrho},$
which concludes the claim. It remains to check the compactness of the above embedding. To prove this, take a bounded sequence $\{v_{k}\}_{k}\subset E$, then one can find a subsequence $\{v_{k_{j}}\}_{j}$ and $v$ in $E$ satisfying
$v_{k_{j}}\rightharpoonup v~\text{in}~E,~v_{k_{j}}\to v~\text{in}~L^{\theta}(B_{R})~~\text{for all}~R>0~~\text{and}~~\theta\geq 1,~v_{k_{j}}\to v~\text{a.e. in}~\mathbb{R}^{n}~\text{as}~j\to\infty,$
we thank to Lemma 2.5. Now, thanks to the Egoroff’s theorem, we can notice that for any open ball $B_{R}$ and $\delta>0$, there exists $A_{\delta}\subset B_{R}$ with $|A_{\delta}|<\delta$ satisfying $v_{k_{j}}\to v$ uniformly in $B_{R}\setminus A_{\delta}$ as $j\to\infty$. Further, let us write
(7.1) $\lim_{j\to\infty}\int_{\mathbb{R}^{n}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$ $=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$
@@
$\quad+\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}\setminus B_{R}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=J_{1}+J_{2}+J_{3}.$
Suppose that $1<t<\frac{n}{\gamma}$ and $1<r<\frac{n}{t\gamma}$ such that $\frac{1}{\tau}+\frac{1}{\tau^{\prime}}=1$ for all $\tau\in\{t,r\}$. Hence, by using the Hölder’s inequality and Lemma 2.5, we have
@@
(7.2) $J_{1}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$
Moreover, we also have
@@
(7.3) $J_{2}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{B_{R}\setminus A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$
@@
(7.4) $J_{3}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}\setminus B_{R}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$
In virtue of (7.2), (7.3) and (7.4), we obtain from (7.1) that $v_{k_{j}}\to v$ in $L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ as $j\to\infty$. This completes the proof. ∎
@@
$v_{k}\rightharpoonup v~\text{in}~E,~v_{k}\to v~\text{in}~L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)~~\text{for all}~~\varrho\geq p^{\ast}~\text{and}~v_{k}\to v~\text{a.e. in}~\mathbb{R}^{n}~\text{as}~k\to\infty,$
we thank to Theorem 1.5. This infers that $\|v\|_{\varrho,\gamma}=\lim_{k\to\infty}\|v_{k}\|_{\varrho,\gamma}=1$. Moreover, due to the weak lower semicontinuity of the norm, we have $\|v\|\leq\liminf_{k\to\infty}\|v_{k}\|=S_{\varrho}\leq\|v\|$. It follows that the best constant $S_{\varrho}$ is achieved. This finishes the proof. ∎
8. Existence of nontrivial weak solutions: Proof of Theorem 1.8
In this section, we shall use the mountain pass theorem, standard topological tools, singular Adams’ inequality, a new concentration-compactness principle due to Lions (see, for example, Theorem 1.1 and Theorem 1.2) and Theorem 1.5 to prove Theorem 1.8. Evidently, the problem ($\mathcal{P}$) has a variational structure. Indeed, since we are interested in studying the problem ($\mathcal{P}$), we define the Euler–Lagrange variational energy functional $J\colon E\to\mathbb{R}$ by
(8.1) $J(u):=\frac{1}{p}\|\Delta u\|^{p}_{p}+\frac{2}{n}\|\Delta u\|^{\frac{n}{2}}_{\frac{n}{2}}-\int_{\mathbb{R}^{n}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x,~\forall~u\in E.$
@@
$\langle{J^{\prime}(u),v}\rangle:=\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u\Delta v\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u\Delta v\,\mathrm{d}x-\int_{\mathbb{R}^{n}}\frac{g(x,u)v}{|x|^{\gamma}}\,\mathrm{d}x,~\forall~u,v\in E.$
Moreover, the critical points of $J$ are exactly weak solutions of the problem ($\mathcal{P}$). To prove our existence results, we need a version of the mountain pass theorem of Rabinowitz [45], as stated below.
@@
(8.2) $\mathcal{I}(u_{k})\to c\quad\text{and}\quad\sup_{\|\varphi\|_{X}=1}|\langle{\mathcal{I}^{\prime}(u_{k}),\varphi}\rangle|\to 0\quad\text{as}\quad k\to\infty$
@@
(The mountain pass theorem) Let $X$ be a real Banach space and $\mathcal{I}\in C^{1}(X,\mathbb{R})$. Suppose $\mathcal{I}(0)=0$ and there hold:
@@
    There exist two constants $\alpha,\rho>0$ such that $\mathcal{I}(u)\geq\alpha$ for all $u\in X$ with $\|u\|=\rho$;
@@
    There exists $e\in X$ satisfying $\|e\|>\rho$ such that $\mathcal{I}(e)<0$.
@@
$\Gamma:=\big\{\gamma\in C^{1}([0,1],X)\colon\gamma(0)=0,~\gamma(1)=e\big\}.$
@@
$c:=\inf_{\gamma\in\Gamma}\max_{t\in[0,1]}\mathcal{I}(\gamma(t))\geq\alpha$
and there exists a $(PS)_{c}$ sequence $\{u_{k}\}_{k}$ for $\mathcal{I}$ in $X$. Consequently, if $\mathcal{I}$ satisfies the Palais-Smale condition, then $\mathcal{I}$ possesses a critical value $c\geq\alpha$.
@@
Let the hypotheses ($\bf g{1}$)($\bf g{4}$) be satisfied. Then the following results are true:
@@
Let $t,t^{\prime}>1$ be such that $\frac{1}{t}+\frac{1}{t^{\prime}}=1$. Moreover, we assume $\kappa>0$ such that $\alpha t^{\prime}\kappa^{\frac{n}{n-2}}\leq\beta_{\gamma,n}$ whenever $\|u\|\leq\kappa$. From (1.15), we obtain by using the Hölder’s inequality, Theorem 1.5 and Corollary 2.6 that
@@
$\leq C(\zeta\|u\|^{\tau}+\|u\|^{q}),~\forall~\zeta>0,$
where we have used Theorem 1.1 to obtain the constant $C>0$, which is given by